# Online Bayesian Experimental Design for Partially Observable Dynamical Systems

### v1

@@ Abstract
Bayesian experimental design (BED) provides a principled framework for optimizing data collection, but existing approaches do not apply to crucial real-world settings such as dynamical systems with partial observability, where only noisy and incomplete observations are available. These systems are naturally modeled as state-space models (SSMs), where latent states mediate the link between parameters and data, making the likelihood—and thus information-theoretic objectives like the expected information gain (EIG)—intractable. In addition, the dynamical nature of the system requires online algorithms that update posterior distributions and select designs sequentially in a computationally efficient manner. We address these challenges by deriving new estimators of the EIG and its gradient that explicitly marginalize latent states, enabling scalable stochastic optimization in nonlinear SSMs. Our approach leverages nested particle filters (NPFs) for efficient online inference with convergence guarantees. Applications to realistic models, such as the susceptible–infected–recovered (SIR) and a moving source location task, show that our framework successfully handles both partial observability and online computation.
@@ 1 INTRODUCTION
Designing experiments that yield maximally informative data is a fundamental problem across science and engineering. The ability to guide data collection efficiently can accelerate learning in diverse domains, including material discovery (lei2021bayesian; lookman2019active), DNA sequencing (weilguny2023dynamic), sensor networks (wu2023offline), and drug discovery (lyu2019ultra). Formally, given a known model structure, let $\boldsymbol{\theta}$ represent the parameters of interest and $\boldsymbol{\xi}$ the design chosen for the experiment, which determines the corresponding observation $\boldsymbol{y}$. The objective is to select designs that maximize the information gained about $\boldsymbol{\theta}$, thereby reducing uncertainty as more data becomes available.
@@
Bayesian experimental design (BED) provides a principled framework for combining Bayesian inference with data acquisition (lindley1956measure; huan2024optimal). In this setting, prior beliefs about the parameters $p(\boldsymbol{\theta})$ are updated as data are collected, yielding a posterior distribution that quantifies uncertainty. This update is driven by the likelihood $p(\boldsymbol{y}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi})$, which links parameters to observations under a given design $\boldsymbol{\xi}$. When this process is repeated adaptively, using past observations to guide future design choices, we obtain the sequential setting known as Bayesian adaptive design (BAD) (Rainforth24). At the experiment $t$, the prior is $p(\boldsymbol{\theta}\!\mid\!h_{t-1})$, where $h_{t-1}=\{\boldsymbol{\xi}_{1:t-1},\boldsymbol{y}_{1:t-1}\}$ is the history of past designs and observations. The next design $\boldsymbol{\xi}_{t}$ is then chosen to maximize some utility function, commonly the expected information gain (EIG), which requires evaluating $p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi}_{t})$.
@@
A common assumption in BED is that the likelihood $p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi}_{t})$ is tractable and available in closed form. While this holds in some settings, it is unrealistic in many real-world scenarios where the system cannot be fully observed. In particular, dynamical systems often involve latent states that evolve over time and are only observed through noisy, partial measurements. Such systems are naturally modeled as state-space models (SSMs) (Sarkka23), where the likelihood is defined only indirectly through integration over the hidden states, and is therefore intractable.
@@
Several BED approaches address implicit likelihoods, i.e., cases where we can simulate observations from the model but cannot evaluate $p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi}_{t})$ in closed form. In principle, such methods could also be applied or extended to SSMs, since they provide a way to simulate samples. Most existing BED approaches rely on amortization, where a surrogate (e.g., a neural network) is trained offline on a fixed dataset to approximate the likelihood, and then reused at deployment. Such methods have been proposed, mostly in static BED settings  (Kleinegesse20; Dehideniya18), with recent extensions to sequential settings via mutual-information bounds (Ivanova21) or density-ratio estimation (Kleinegesse21).
In contrast, we focus on the non-amortized setting: designs are optimized sequentially, and their accuracy can be improved arbitrarily by allocating more computational resources at each step. This setting is particularly relevant in applications where experiments themselves are costly, such as adaptive clinical trials (chaloner1995bayesian) and astronomical observations (loredo2004bayesian). In such cases, the cost of additional computation between measurements is negligible compared to the cost of data collection. Our aim is therefore to develop methods that retain full sequential adaptability while ensuring that design optimization at each step remains as accurate as resources allow. We provide further discussion of related work in Section 4.
@@
Contributions. In this paper, we propose a novel framework for BAD in partially observable dynamical systems, where likelihoods are intractable and online inference is required. Our main contributions are (i) new estimators of EIG and its gradient, specifically derived for state-space models with latent states, enabling design optimization under partial observability, and (ii) an online BAD approach based on nested particle filters (NPFs) (Crisan18bernoulli), which recursively approximates the joint posterior over parameters and latent states without reprocessing the full history. In representative experiments, we show that our framework successfully enables BED under partial observability and online constraints.
2 PROBLEM SETTING
2.1 Bayesian adaptive design (BAD)
The goal of BAD is to select, at each experiment $t$, a design $\boldsymbol{\xi}_{t}$ that maximizes the EIG about the parameters:
$\boldsymbol{\xi}_{t}^{\star}=\arg\max_{\boldsymbol{\xi}_{t}\in\Omega}\;\mathcal{I}(\boldsymbol{\xi}_{t}),$
where $\Omega$ is the design space.
The EIG quantifies the reduction in posterior uncertainty about $\boldsymbol{\theta}$ after observing $\boldsymbol{y}_{t}$ under design $\boldsymbol{\xi}_{t}$:
$\mathcal{I}$ $(\boldsymbol{\xi}_{t})=\mathbb{E}_{p(\boldsymbol{y}_{t}\mid\boldsymbol{\xi}_{t})}\!\bigg[\,\mathcal{H}[p(\boldsymbol{\theta}\!\mid\!h_{t-1})]-\mathcal{H}[p(\boldsymbol{\theta}\!\mid\!h_{t})]\bigg]$
$=\mathbb{E}_{p(\boldsymbol{\theta}\mid h_{t-1})\,p(\boldsymbol{y}_{t}\mid\boldsymbol{\theta},\boldsymbol{\xi}_{t})}\!\big[\log p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi}_{t})-\log p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\xi}_{t})\big],$ (1)
@@ 2.1 Bayesian adaptive design (BAD)
with $h_{t-1}=(\boldsymbol{\xi}_{1:t-1},\boldsymbol{y}_{1:t-1})$ the history, $p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi}_{t})$ the likelihood, $p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\xi}_{t})$ the marginal predictive distribution, and $\mathcal{H}$ denoting entropy.
@@
In many applications, however, observations $\boldsymbol{y}_{t}$ are not generated directly by $\boldsymbol{\theta}$, but through an underlying dynamical system with latent states $\boldsymbol{x}_{0:t}$. Such cases are naturally modeled as SSMs, resulting in a setting where the likelihood is not available in closed form, rendering standard BAD methods inapplicable.
@@ 2.3 Particle filters (PFs)
In SSMs, the likelihood $p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi}_{t},h_{t-1})$ is defined only through marginalization over latent states, which is intractable in general. Particle filters (PFs), or sequential Monte Carlo methods (Gordon93; Doucet00; Doucet01; Djuric03), provide a standard solution by recursively approximating the filtering or posterior distribution, by representing it as a weighted set of $M$ particles,
@@
At each step, particles $\boldsymbol{x}_{t}^{(m)}$ are propagated forward using the state transition $f$, while their weights $w_{\boldsymbol{x},t}^{(m)}$ are updated in proportion to the observation model $g$. This allows the algorithm to adaptively focus on regions of high likelihood and provides a flexible representation of complex, nonlinear distributions.
Standard PFs assume that both the parameters $\boldsymbol{\theta}$ and the designs $\boldsymbol{\xi}_{1:t}$ are known and fixed, so that only the latent states need to be inferred. In BAD, however, the parameters $\boldsymbol{\theta}$ are also unknown and must be estimated. This motivates the use of NPFs, which extend the PF framework to incorporate parameter inference.
@@ 2.4 Nested particle filters (NPFs)
NPFs (Crisan18bernoulli) extend PFs to jointly estimate states and parameters in an online manner, targeting the joint posterior $p(\mathrm{d}\boldsymbol{x}_{0:t},\mathrm{d}\boldsymbol{\theta}\!\mid\!\boldsymbol{y}_{1:t},\boldsymbol{\xi}_{1:t})$. An NPF employs two intertwined layers of PFs: (i) an inner layer that estimates the state distribution conditional on each parameter particle
@@
This nested structure is related to sequential Monte Carlo squared (SMC2) (chopin2013smc2), but with a key computational improvement: instead of re-running inner filters from scratch, NPFs perturb parameters using a jittering kernel,
$\boldsymbol{\theta}_{t}^{(m)}\sim\kappa_{N}(\mathrm{d}\boldsymbol{\theta}\!\mid\!\boldsymbol{\theta}_{t-1}^{(m)}),$
so that the inner filters are updated a single step forward. This avoids reprocessing past data and yields linear cost $\mathcal{O}(t)$, compared to the quadratic $\mathcal{O}(t^{2})$ cost of SMC2.
Under suitable conditions, jittering prevents particle degeneracy and ensures consistency, providing asymptotic convergence guarantees (Crisan18bernoulli). This has motivated several extensions and further analysis, including Gaussian approximations (Perez-Vieites21) and Rao–Blackwellized variants (Fang23).
3 PROPOSED FRAMEWORK
We propose a new BAD framework for partially observable dynamical systems, deriving recursive estimators of the EIG and its gradient. This enables online optimization in continuous design spaces, $\boldsymbol{\xi}_{t}\in\mathbb{R}^{d_{\xi}}$, using stochastic gradient ascent (SGA). We further establish consistency of the proposed EIG estimator, and conclude the section by presenting the overall algorithmic scheme required for practical implementation.
@@ 3.1 EIG in SSMs
The EIG expression in (2.1) shows the dependency on the likelihood $p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi}_{t})$ and the marginal likelihood (evidence) $p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\xi}_{t})=\mathbb{E}_{p(\boldsymbol{\theta}\mid h_{t-1})}[\,p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi}_{t})]$. Even when the likelihood is available in closed-form, estimating $\mathcal{I}(\boldsymbol{\xi}_{t})$ is doubly intractable because the evidence appears inside the outer expectation that defines the utility, typically needing nested Monte Carlo (NMC) methods or other approximations (Rainforth18; Foster19).
@@
respectively. Substituting these into (2.1) yields
@@ 3.2 Gradient of the EIG
Using Fisher’s identity (Douc14) on the EIG of (4), we obtain a gradient representation that separates (i) derivatives of the likelihood and evidence and (ii) the design score at time $t$. Let
@@
$~+\log\!\left[\,\frac{L_{\boldsymbol{\theta},\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})}{Z_{\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})}\right]s_{\boldsymbol{\theta},\boldsymbol{\xi}_{t}}(\boldsymbol{x}_{t-1:t},\boldsymbol{y}_{t})\Bigg]$ (5)
@@
with any of the terms vanishing when they do not depend on $\boldsymbol{\xi}_{t}$. See details in Appendix LABEL:supp-ap:gradient.
@@
In the next subsection, we show how this quadratic bottleneck can be avoided using the NPF framework, yielding Monte Carlo gradient estimators that require only a single forward propagation step per iteration, i.e., with complexity $\mathcal{O}(t)$.

### v2

# Online Bayesian Experimental Design for Partially Observed Dynamical Systems

@@ Abstract
Bayesian experimental design (BED) provides a principled framework for optimizing data collection by choosing experiments that are maximally informative about unknown parameters. However, existing methods cannot deal with the joint challenge of (a) partially observable dynamical systems, where only noisy and incomplete observations are available, and (b) fully online inference, which updates posterior distributions and selects designs sequentially in a computationally efficient manner. Under partial observability, dynamical systems are naturally modeled as state-space models (SSMs), where latent states mediate the link between parameters and data, making the likelihood—and thus information-theoretic objectives like the expected information gain (EIG)—intractable. We address these challenges by deriving new estimators of the EIG and its gradient that explicitly marginalize latent states, enabling scalable stochastic optimization in nonlinear SSMs. Our approach leverages nested particle filters for efficient online state-parameter inference with convergence guarantees. Applications to realistic models, such as the susceptible–infectious–recovered (SIR) and a moving source location task, show that our method successfully handles both partial observability and online inference.
Bayesian experimental design, Bayesian inference, Monte Carlo
@@ 1 Introduction
Designing experiments that yield maximally informative data is a fundamental problem across science and engineering. The ability to guide data collection efficiently can accelerate learning in diverse domains, including material discovery (Lei et al., 2021; Lookman et al., 2019), DNA sequencing (Weilguny et al., 2023), sensor networks (Wu et al., 2023), and drug discovery (Lyu et al., 2019). Formally, given a known model structure, let $\boldsymbol{\theta}$ represent the parameters of interest and $\boldsymbol{\xi}$ the design chosen for the experiment, which determines the corresponding observation $\boldsymbol{y}$. The objective is to select designs that maximize the information gained about $\boldsymbol{\theta}$, thereby reducing uncertainty as more data becomes available.
@@
Bayesian experimental design (BED) provides a principled framework for combining Bayesian inference with data acquisition (Lindley, 1956; Huan et al., 2024). In this setting, prior beliefs about the parameters $p(\boldsymbol{\theta})$ are updated as data are collected, yielding a posterior distribution that quantifies uncertainty. This update is driven by the likelihood $p(\boldsymbol{y}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi})$, which links parameters to observations under a given design $\boldsymbol{\xi}$. When this process is repeated adaptively, using past observations to guide future design choices, we obtain the sequential BED setting (Rainforth et al., 2024). At the experiment $t$, the prior is $p(\boldsymbol{\theta}\!\mid\!h_{t-1})$, where $h_{t-1}=\{\boldsymbol{\xi}_{1:t-1},\boldsymbol{y}_{1:t-1}\}$ is the history of past designs and observations. The next design $\boldsymbol{\xi}_{t}$ is then chosen to maximize some utility function, commonly the expected information gain (EIG), which requires evaluating $p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi}_{t})$.
@@
A common assumption in sequential BED is that the likelihood $p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi}_{t})$ is tractable and available in closed form. While this holds in some settings, it is unrealistic in many real-world scenarios where the system cannot be fully observed. In particular, dynamical systems often involve latent states that evolve over time and are only observed through noisy, partial measurements. Such systems are naturally modeled as state-space models (SSMs) (Särkkä and Svensson, 2023), where the likelihood is defined only indirectly through integration over the hidden states, and is therefore intractable.
@@
Several BED approaches address implicit likelihoods, i.e., cases where we can simulate observations from the model but cannot evaluate $p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi}_{t})$ in closed form. In principle, such methods could also be applied or extended to SSMs, since they provide a way to simulate samples. Most existing implicit BED approaches rely on amortization, where a surrogate (e.g., a neural network) is trained offline on a fixed dataset to approximate the likelihood, and then reused at deployment. Such methods have been proposed, mostly in static BED settings  (Kleinegesse and Gutmann, 2020; Dehideniya et al., 2018), with recent extensions to sequential settings via mutual-information bounds (Ivanova et al., 2021) or density-ratio estimation (Kleinegesse and Gutmann, 2021).
In contrast, we focus on a non-amortized, fully online regime: designs are optimized sequentially using the current posterior and without offline training. This regime is motivated by applications where (i) decisions are made over long horizons, (ii) the system is only partially observed, and (iii) computational budget between measurements can vary—often prioritizing reliability over ultra-fast deployment, e.g., adaptive clinical trials (Chaloner and Verdinelli, 1995) and astronomical observations (Loredo, 2004). A key practical issue in such settings is long-horizon feasibility. In partially observed dynamical models, robust inference/design schemes that repeatedly rely on the full data history and long latent trajectories can incur computation and memory costs that grow with $T$, which may become prohibitive in practice. Our aim is therefore to develop methods that keep inference and design updates online (with linear cost in $T$ and constant memory in the horizon length) while retaining convergence guarantees for the estimators used in design optimization. We discuss these trade-offs and connections to prior work in Section 4.
@@
Contributions. In this paper, we propose a novel method for sequential BED in partially observable dynamical systems, where likelihoods are intractable and online inference is required. Our main contributions are (i) new estimators of EIG and its gradient, specifically derived for state-space models with latent states, enabling design optimization under partial observability; (ii) an online sequential BED approach that couples these estimators with nested particle filters (NPFs, Crisan and Míguez, 2018), which recursively updates the joint parameter–state posterior without reprocessing the full history, yielding inference and design updates with linear cost in the horizon; and (iii) asymptotic consistency guarantees for the resulting EIG estimators when implemented with our approach.
2 Background
2.1 Sequential Bayesian Experimental Design
The goal of sequential BED is to select, at each experiment $t$, a design $\boldsymbol{\xi}_{t}$ that maximizes the expected information gain (EIG, Lindley, 1956) about the parameters: $\boldsymbol{\xi}_{t}^{\star}=\arg\max_{\boldsymbol{\xi}_{t}\in\Omega}\;\mathcal{I}(\boldsymbol{\xi}_{t})$, where $\Omega$ is the design space. The EIG quantifies the reduction in posterior uncertainty about $\boldsymbol{\theta}$ after observing $\boldsymbol{y}_{t}$ under design $\boldsymbol{\xi}_{t}$, and is defined as:
$\mathcal{I}$ $(\boldsymbol{\xi}_{t})=\mathbb{E}_{p(\boldsymbol{y}_{t}\mid\boldsymbol{\xi}_{t})}\!\bigg[\,\mathcal{H}[p(\boldsymbol{\theta}\!\mid\!h_{t-1})]-\mathcal{H}[p(\boldsymbol{\theta}\!\mid\!h_{t})]\bigg]$ (1)
$=\mathbb{E}_{p(\boldsymbol{\theta}\mid h_{t-1})\,p(\boldsymbol{y}_{t}\mid\boldsymbol{\theta},\boldsymbol{\xi}_{t})}\!\big[\log p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi}_{t})-\log p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\xi}_{t})\big],$
@@ 2.1 Sequential Bayesian Experimental Design
with $h_{t-1}\!=\!\{\boldsymbol{\xi}_{1:t-1},\boldsymbol{y}_{1:t-1}\}$ the history, $p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi}_{t})$ the likelihood, $p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\xi}_{t})$ the marginal predictive distribution, and $\mathcal{H}$ denoting entropy.
@@
In many applications, however, observations $\boldsymbol{y}_{t}$ are not generated directly by $\boldsymbol{\theta}$, but through an underlying dynamical system with latent states $\boldsymbol{x}_{0:t}$. Such cases are naturally modeled as SSMs, for which the likelihood is not available in closed form, rendering standard BED methods inapplicable.
@@ 2.3 Particle Filters (PFs)
In SSMs, the likelihood $p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi}_{t},h_{t-1})$ typically has no closed form, as it requires marginalising the latent state,
$p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\!\boldsymbol{\xi}_{t},\!h_{t-1})\!=\!\int\!g(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{x}_{t},\boldsymbol{\theta},\boldsymbol{\xi}_{t})\,p(\mathrm{d}\boldsymbol{x}_{t}\!\mid\!\boldsymbol{y}_{1:t},\boldsymbol{\theta},\boldsymbol{\xi}_{1:t}).$ (4)
Particle filters (PFs, Gordon et al., 1993; Doucet et al., 2000, 2001; Djurić et al., 2003) provide a standard solution by recursively approximating the filtering or posterior distribution with a weighted set of $M$ particles,
@@
At each time step, particles $\boldsymbol{x}_{t}^{(m)}$ are propagated forward using the state transition $f$, while their weights $w_{\boldsymbol{x},t}^{(m)}$ are updated in proportion to the observation model $g$. This allows the algorithm to adaptively focus on regions of high likelihood and provides a flexible representation of complex, nonlinear distributions.
Standard PFs assume that both the parameters $\boldsymbol{\theta}$ and the designs $\boldsymbol{\xi}_{1:t}$ are known and fixed, so that only the latent states need to be inferred. In sequential BED, however, the parameters $\boldsymbol{\theta}$ are also unknown and must be estimated. This motivates the use of NPFs, which extend the PF framework to incorporate parameter inference.
@@ 2.4 Nested Particle Filters (NPFs)
NPFs (Crisan and Míguez, 2018) extend PFs to jointly estimate states and parameters online, targeting the joint posterior $p(\mathrm{d}\boldsymbol{x}_{0:t},\mathrm{d}\boldsymbol{\theta}\!\mid\!\boldsymbol{y}_{1:t},\boldsymbol{\xi}_{1:t})$. An NPF employs two intertwined layers of PFs: (i) an inner layer that estimates the state distribution conditional on each parameter particle
@@
The parameter weights are updated using an estimate of the likelihood in (4) obtained from the inner filters.
This nested structure is related to sequential Monte Carlo squared (SMC2, Chopin et al., 2013), but differs in a key aspect. In SMC2, when parameters are moved (e.g., via MCMC rejuvenation), the inner filters are typically re-run using the full data history, leading to a cost that grows quadratically with time. This becomes problematic in long-horizon settings, both in runtime and memory. NPFs instead use a jittering kernel to perturb parameter particles,
$\boldsymbol{\theta}_{t}^{(m)}\sim\kappa_{M}\big(\mathrm{d}\boldsymbol{\theta}\!\mid\!\boldsymbol{\theta}_{t-1}^{(m)}\big),$
with a variance chosen to satisfy the regularity conditions required for the NPF convergence results (Crisan and Míguez, 2018), and tuned in practice to balance exploration and stability. This controlled perturbation enables a purely online update: each inner PF is advanced a single step (rather than reprocessing past data), yielding linear cost in $t$.
Several extensions replace the inner PF with more structured approximations (e.g., Gaussian or Rao–Blackwellised variants) while retaining theoretical guarantees under modified assumptions (Pérez-Vieites and Míguez, 2021; Fang et al., 2023).
3 Proposed Method
We propose a new sequential BED method for partially observable dynamical systems, deriving recursive estimators of the EIG and its gradient. This enables online optimization in continuous design spaces, $\boldsymbol{\xi}_{t}\!\in\!\mathbb{R}^{d_{\xi}}$, using stochastic gradient ascent (SGA). We further establish consistency of the proposed EIG estimator, and conclude the section by presenting the overall algorithmic scheme.
@@ 3.1 EIG in SSMs
The EIG expression in (1) shows the dependency on the likelihood $p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi}_{t})$ and the marginal likelihood (evidence) $p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\xi}_{t})=\mathbb{E}_{p(\boldsymbol{\theta}\mid h_{t-1})}[\,p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi}_{t})]$. Even when the likelihood is available in closed-form, estimating $\mathcal{I}(\boldsymbol{\xi}_{t})$ is doubly intractable because the evidence appears inside the outer expectation that defines the utility, typically needing nested Monte Carlo (NMC) methods or other approximations (Rainforth et al., 2018; Foster et al., 2019).
@@
respectively. Substituting these into (1) yields
@@ 3.2 Gradient of the EIG
Using Fisher’s identity (Douc et al., 2014) on the EIG of (5), we obtain a gradient representation that separates (i) derivatives of the likelihood and evidence and (ii) the design score at time $t$. Let
@@
$~+\log\!\left[\,\frac{L_{\boldsymbol{\theta},\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})}{Z_{\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})}\right]s_{\boldsymbol{\theta},\boldsymbol{\xi}_{t}}(\boldsymbol{x}_{t-1:t},\boldsymbol{y}_{t})\Bigg],$ (6)
@@
with any of the terms vanishing when they do not depend on $\boldsymbol{\xi}_{t}$. See details in Appendix A.
@@
In the next subsection, we show how this quadratic bottleneck can be avoided using NPFs, yielding Monte Carlo gradient estimators that require only a single forward propagation step per iteration, i.e., with complexity $\mathcal{O}(t)$.